# CHIPS: Efficient CLIP Adaptation via Curvature-aware Hybrid Influence-based Data Selection

### v1

@@ 2.4 Learnability and Domain Relevance
Let $\tilde{q}\propto A(\cdot)\,w_{\mathrm{L}}(\cdot)$ and $q\propto w_{\mathrm{R}}(\cdot)\,\tilde{q}$ be the selection distributions with/without $w_{\mathrm{R}}$. Because $w_{\mathrm{R}}\in[a,b]$ with $a=\sigma(-1),\,b=\sigma(1)$, the density ratio $r(z)=\frac{q(z)}{\tilde{q}(z)}=\frac{w_{\mathrm{R}}(z)}{\mathbb{E}_{\tilde{q}}[w_{\mathrm{R}}]}$ satisfies $r(z)\in[a/b,\,b/a]=[e^{-1},e]$, which implies a bounded selection drift $D_{\mathrm{KL}}(q\,\|\,\tilde{q})\leq 1$ nat. Thus $w_{\mathrm{R}}$ also limits how far the selection moves from the base alignment-driven distribution, mitigating catastrophic forgetting of general-domain knowledge. We analyze $\beta$ in Sec. 4.2 to balance adaptation and retention.
@@ 3.2 Main Results
As shown in Tab. 2, across all retention ratios, CHIPS achieves the best average performance on 17 medical tasks among data selection methods, with Medical Avg scores of 27.03, 28.20, and 29.96, exceeding the second–best method by +0.57, +1.57, and +3.68 points, respectively. Notably, with only 10% of the full BIOMEDICA, CHIPS outperforms a 50% random subset (27.03 vs. 26.26); with 30%, it achieves 95.1% of the full-dataset performance (29.96 vs. 31.51). It is also competitive with specialized medical CLIP models pre-trained on larger datasets: at $r=30\%$, CHIPS slightly surpasses BMCLIP (29.96 vs. 29.86) and consistently outperforms PubMedCLIP and BioMedCLIP across $r$. Even though CHIPS is not uniformly best on every specialty, it yields the highest overall average. Importantly, CHIPS better preserves general-domain abilities than the previous SOTA selection method TracIn: for CLS, CHIPS retains 90.1%, 89.3%, and 87.2% of the vanilla model at $r=10\%,20\%,30\%$ (vs. 88.9%, 88.0%, 87.0% for TracIn), and for RET it is comparable, slightly lower at $r=10\%$ but higher at $r\geq 20\%$, yielding a higher average RET across retention ratios. Overall, CHIPS delivers the strongest medical performance while incurring less degradation on the general domain than TracIn.
@@ 4.1 Alignment
In Eq. (8), $\alpha$ linearly mixes positive and negative pair curvature to form $\mathbf{M}$, which reweights gradients in $A(z)=\mathbf{g}_{\vartheta}(z)^{\top}\mathbf{M}^{-1}\mathbf{u}_{\vartheta}$. As shown in Fig. 2(B), Medical CLS peaks at $\alpha{=}0.6$ (27.05) and remains essentially tied at $0.8$ (27.03), while both smaller ($0.2$: 25.92; $0.4$: 25.66) and larger ($1.0$: 26.07) values underperform by about 1–1.4 points. General CLS is almost flat across $\alpha$ (48.36–48.46), with a modest dip at $0.6$ (47.88), indicating that $\alpha$ mainly governs target-domain discrimination with negligible general-domain impact. Mechanistically, moderate $\alpha$ balances attraction to positives and repulsion from negatives, improving the conditioning of $\mathbf{M}$ and yielding a more discriminative alignment direction; extremes overweight one term and reduce $A(z)$’s utility. We therefore recommend $\alpha\in[0.6,0.8]$ (default $0.6$ for medical focus; $0.4$ if slightly prioritizing general-domain retention).
@@ G.1 Numerical Totals
We instantiate Eqs. (25)–(34) with the fixed values above. We use $E=10$ epochs for TracIn and $I=5$ CG iterations for TRAK and CHIPS.

### v2

@@ 2.4 Learnability and Domain Relevance
Hence, $w_{\mathrm{R}}$ acts as a gentle target-domain bias rather than a hard filter: it nudges selection toward target-relevant samples while preventing the final distribution from drifting too far from the original alignment-driven one. This helps balance adaptation and retention, mitigating catastrophic forgetting of general-domain knowledge.
@@ 3.2 Main Results
As shown in Tab. 2, across all retention ratios, CHIPS achieves the best average performance on 17 medical tasks among data selection methods, with Medical Avg scores of 27.03, 28.20, and 29.96, exceeding the second-best method by +0.57, +1.57, and +3.68 points, respectively. Notably, with only 10% of the full BIOMEDICA, CHIPS outperforms a 50% random subset (27.03 vs. 26.26); with 30%, it achieves 95.1% of the full-dataset performance (29.96 vs. 31.51). It is also competitive with specialized medical CLIP models pre-trained on larger datasets: at $r=30\%$, CHIPS slightly surpasses BMCLIP (29.96 vs. 29.86) and consistently outperforms PubMedCLIP and BioMedCLIP across $r$. Even though CHIPS is not uniformly best on every specialty, it yields the highest overall average. Importantly, CHIPS better preserves general-domain abilities than the previous SOTA selection method TracIn: for CLS, CHIPS retains 90.1%, 89.3%, and 87.2% of the vanilla model at $r=10\%,20\%,30\%$ (vs. 88.9%, 88.0%, 87.0% for TracIn), and for RET it is comparable, slightly lower at $r=10\%$ but higher at $r\geq 20\%$, yielding a higher average RET across retention ratios. Overall, CHIPS delivers the strongest medical performance while incurring less degradation on the general domain than TracIn.
@@ 4.1 Alignment
In Eq. (7), $\alpha$ linearly mixes positive and negative pair curvature to form $\mathbf{M}$, which reweights gradients in $A(z)=\mathbf{g}_{\vartheta}(z)^{\top}\mathbf{M}^{-1}\mathbf{u}_{\vartheta}$. As shown in Fig. 2(B), Medical CLS peaks at $\alpha{=}0.6$ (27.05) and remains essentially tied at $0.8$ (27.03), while both smaller ($0.2$: 25.92; $0.4$: 25.66) and larger ($1.0$: 26.07) values underperform by about 1-1.4 points. General CLS is almost flat across $\alpha$ (48.36-48.46), with a modest dip at $0.6$ (47.88), indicating that $\alpha$ mainly governs target-domain discrimination with negligible general-domain impact. Mechanistically, moderate $\alpha$ balances attraction to positives and repulsion from negatives, improving the conditioning of $\mathbf{M}$ and yielding a more discriminative alignment direction; extremes overweight one term and reduce $A(z)$’s utility. We therefore recommend $\alpha\in[0.6,0.8]$ (default $0.6$ for medical focus; $0.4$ if slightly prioritizing general-domain retention).
@@ G.1 Numerical Totals
We instantiate Eqs. (24)-(33) with the fixed values above. We use $E=10$ epochs for TracIn and $I=5$ CG iterations for TRAK and CHIPS.